OpenStudy (anonymous):

The width of Ben's rectangular family room is 3 ft. less than the length. The perimeter is 70 ft. Which equation can be used to find the length, l, of the room?

jhonyy9 (jhonyy9):

w=l-3 ft P=70 ft P = 2l +2w P = 2l +2(l-3) P = 2l +2l -6 P = 4l -6 70 = 4l -6 76 = 4l l = 76/4 l = 19 ft